OpenStudy (anonymous):

Solve 2x + 5y = 10 for y. y = 2x + 2 y = –2x + 10 y = -2/5x + 2 y = 2/5x + 10

OpenStudy (anonymous):

5y= -2x+10 y= -2/5 x+10 So it would be D

OpenStudy (lopezking1):

Firstly move the 2x to the other side - via subtraction - so the y part is isolated: 5y = 10 - 2x Then move the 5 over to the other side - via division - so the y stands alone: y = (10 - 2x) / 5 Finally divide through - simplify: y = 2 - 2/5x Which when rearranged is answer 3: y = -2/5 x + 2 C

OpenStudy (luigi0210):

Nope

OpenStudy (luigi0210):

Warrior messed up

OpenStudy (anonymous):

My bad. It would be C

OpenStudy (anonymous):

i thought you were done!

OpenStudy (anonymous):

it's c??

OpenStudy (lopezking1):

@pamelasaucedo

OpenStudy (lopezking1):

yea

OpenStudy (luigi0210):

good job